Find four consecutive multiples of five such that twice the sum of the two biggest exceeds twice the smallest by ten.
Answer by Edwin McCravy(20054)

Find four consecutive multiples of five such that twice the sum of the two biggest exceeds twice the smallest by ten.

5k     = smallest multiple of 5
5(k+1) = next to smallest multiple of 5
5(k+2) = next to biggest multiple of 5
5(k+3) = biggest multiple of 5

>>twice the sum of the two biggest... 

sum of the two biggest = 5(k+2) + 5(k+3)

twice the sum of the two biggest = 2[5(k+2) + 5(k+3)]

...exceeds twice the smallest by ten.<<

twice the smallest = 2(5k)

2[5(k+2) + 5(k+3)] = 2(5k) + 10

Can you solve that for k?  If not post again asking how.

Solution: k = -4

smallest multiple of 5         = 5k     = 5(-4)   =         -20
next to smallest multiple of 5 = 5(k+1) = 5(-4+1) = 5(-3) = -15
next to biggest multiple of 5  = 5(k+2) = 5(-4+2) = 5(-2) = -10
biggest multiple of 5          = 5(k+3) = 5(-4+3) = 5(-1) =  -5

So they are -20, -15, -10, -5.    

Edwin